OpenStudy (anonymous):

help? medal will be given * The value of y varies directly with x. What is k the constant of variation for this function, and what is y for the given value of x? y = -1/2 when x=6. Find y when x=-3

OpenStudy (anonymous):

k = 1/12; -4 k = 12; -4 k = –12; 4 k = -1/12; 1/4

hartnn (hartnn):

for direct variation, the equation is y=kx plug in y=-1/2 , x=6 in this and get the value of k. can you ?

OpenStudy (anonymous):

k = -1/12 ?

hartnn (hartnn):

yes! now you have y= (-1/12)x so, just plug in x=-3 and get the value of y.

OpenStudy (anonymous):

y = 1 / 4 ?

hartnn (hartnn):

very good! its correct \[\checkmark\]